Suppose you invest $1500 into an account that looses 4.75% annually. Find your balance after 25 years if compound semi-annually.
Answer by rfer(16322)

A=1500(1+0.0475/2)^2*25
A=1500(1.02375)^50
A=1500(3.23367)
A=$4850.50